Two inlet pipe A, B filled the tank in 15 and 12 hr respectively Outlet pipe C can empty the tank in 4 hrs. If the pipe are opened in order 8,9, and 11 am then at what time tank will be empty.

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
In one hour, A will fill 1/15 of tankB will fill 1/12 of tank C will drain 1/4 of tank.Now assume t=0 at 8amAnd at t it drains. t*(1/15)+ (t-1)*(1/12)- (t-3)*(1/4) =0 Multiply by 60 both side4t+ 5(t-1)- 15(t-3)=0-11t+40=0Time 3.63 hrThat is 3 hr 38 min from 8 am so time will be 11:38 am